OpenStudy (anonymous):

The half-life of a radioactive isotope is the time it takes for a quantity of the isotope to be reduced to half its initial mass. Starting with 150 grams of a radioactive isotope, how much will be left after 3 half lives? round your answer to the nearest gram.

OpenStudy (anonymous):

Start with 150, divide that by two divide that number by two divide that number by two

OpenStudy (ribhu):

after ist half life remaining would be 75gm

OpenStudy (ribhu):

second half life remaining would be 75-37.5 = 37.5 gm.

OpenStudy (ribhu):

after third half life 37.5-37.5/2 =37.5/2 = 18.75 gms

OpenStudy (ribhu):

18.75 gm of isotope would be remaining

OpenStudy (anonymous):

Thank you so much :) may I ask you where you got the 2 from? when you divide 150/2 just so I know what to do for next time.

OpenStudy (ribhu):